OpenStudy (anonymous):

What is the kinetic energy (in eV) of a Broglie wavelength of 1.4x10^-10m. Knowing the mass of a neutron is 1.67x10^-27 kg.

OpenStudy (michele_laino):

please, note that the kinetic energy E of a not-relativistic neutron whose mass is M, is equals to: \[E=\frac{ p ^{2} }{ 2M }\] where p is the momentum of our neutron. By De Broglie relationship, we can write: \[\lambda=\frac{ h }{ p }\] where lambda is the wave length of our neutron, and h is the Planck constant, namely h=6.62*10^-27 erg*sec

OpenStudy (michele_laino):

so after substitution into relationship for E, we get: \[E=\frac{ h ^{2} }{ 2M \lambda ^{2} }\]

OpenStudy (michele_laino):

now, we have to insert your numerical data into equation above, in order to get the kinetic energy E, so we have:

OpenStudy (michele_laino):

\[E \approx 6*10^{-14} erg\]

OpenStudy (michele_laino):

or \[E \approx 6*10^{-21} Joule\] now, keep in mind that \[1eV=1.6*10^{-19 }Joule\] we get: \[E \approx 3.2*10^{-2}eV\]